OpenStudy (anonymous):

Non-Homogeneous equation,i needed help figuring the particular solution to a problem or how to make the guess for it

OpenStudy (anonymous):

y′′ − 3y′ + 2y  =  4x + e^(2x)

OpenStudy (anonymous):

i figured out the complimentary part just need help with particular solution because im coming up empty handed

OpenStudy (anonymous):

in case you are wonderin i made the guess Ax+B+C*e^(2x) also hi Zarkon =)

OpenStudy (zarkon):

what is the solution to \[x^2-3x+2=0\]

OpenStudy (anonymous):

(x-2)(x-1)=0 x=2 x=1 y=c1*e^(2x)+c2*e^(x)

OpenStudy (zarkon):

so x=2 is a solution...that should help you with you guess

OpenStudy (anonymous):

for particular solution? sorry im just learning this but i dont see how

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

can you explain how i should use my complimenary solution root to guess particular solution

OpenStudy (anonymous):

since it is this form 4x + e^(2x) i thought the guess should be in the form Ax+B+C*e^(2x)

OpenStudy (zarkon):

Ax+B+C*x*e^(2x)

OpenStudy (anonymous):

what prompted u to guess the extra x?

OpenStudy (zarkon):

you multiply by x since 2 is a solution to the auxiliary equation

OpenStudy (anonymous):

so if there is more than one root to the auxilary equation? or just 2 in particular

OpenStudy (zarkon):

the 2 since on the RHS of the eaution we have \(e^{2x}\) ... a 2 in the exponent

OpenStudy (zarkon):

equation

OpenStudy (anonymous):

OH My god thankyou thats actually something never explained

OpenStudy (anonymous):

so if your root exists in the exponential equation you should multiply by x to make it equivalent

OpenStudy (zarkon):

if the 2 had been a repeated root then you would have multiplied bt \(x^2\) and not just \(x\)

OpenStudy (anonymous):

yes that makes so much more sense also one last question if there was (x-2)(x-3) but the exponent to was to the power e^(4x) we wouldnt multiply by x right?

OpenStudy (zarkon):

what book do you use

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

for engineers book one sec lemme google it

OpenStudy (anonymous):

but i think a newer version but i cant find it lol

OpenStudy (anonymous):

may your drule empire be everlasting king zarkon

OpenStudy (zarkon):

look at the book page 186

OpenStudy (anonymous):

its loading..............

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

should i post another question in another place or is it ok if i post here

OpenStudy (anonymous):

its a different approach to the standard homogeneous equations so it kinda threw me off

OpenStudy (zarkon):

I have to go away...so make a new post

OpenStudy (anonymous):

KK THANKS AGAIN LORD ZARKON

OpenStudy (zarkon):

np